A window consists of a rectangle surmounted by a semicircle. the rectanglular portion has a width of 2m and a height of 1m.Determine the expression for the exact peremetre of the window
Answer by addingup(3677)

perimeter for your window is 2w+2h+πr (normally the circumference would be π2r but here we have only one-half of the circumference).
-----------------------------------------------
2(2)+2(1) = 6 this is the perimeter of the rectangle.
The semi circle on top of the rectangle:
Width of window: 2 this is the diameter of your semicircle. The radius is one-half: 2/2 = 1
π*1 = perimeter of the half circle = 3.14
Total perimeter:
6+3.14 = 9.14